OpenStudy (anonymous):

For witch operation is the set {0,1} closed? Choose ALL answers that are correct. I will fan and medal! please help!

OpenStudy (freckles):

Do you know what closed under an operation means?

OpenStudy (anonymous):

nope

OpenStudy (freckles):

Closed means if you take any two elements from that and add them you will not get something out that set as an output

OpenStudy (freckles):

For example what you took the element 1 from that set what is 1+1?

OpenStudy (anonymous):

2 lol

OpenStudy (freckles):

Is 2 an element in the set {0,1}?

OpenStudy (anonymous):

no?

OpenStudy (freckles):

Right only 0 and 1 are elements of the set {0,1} 2 isn't an element of the set {0,1} Which means since we tried to add that the set {0,1} is not closed under addition

OpenStudy (freckles):

Try multiplication

OpenStudy (freckles):

Is the set {0,1} closed under multiplication

OpenStudy (freckles):

Try all combinations... 0*0=? 0*1=? 1*0=? 1*1=?

OpenStudy (anonymous):

0 0 0 1

OpenStudy (freckles):

aren't all of those outputs elements of the set {0,1}?

OpenStudy (anonymous):

yea

OpenStudy (freckles):

so it the set {0,1} closed under multiplication?

OpenStudy (anonymous):

yes

OpenStudy (freckles):

yep yep

OpenStudy (freckles):

Do you want to try any other operations?

OpenStudy (anonymous):

yeah sense it is u know all answers that are correct lmao

OpenStudy (freckles):

Ok well you know what closed means so try other operations

OpenStudy (freckles):

by the way you can define new operations but i don't think your teacher is asking you to do that

OpenStudy (freckles):

i think you are to try the operations, +,-,x,/